QUESTION 6 The Hi-Tek Company just decided to save $28,000 a month for the next five years as a safety net for recessionary periods. The money will be set aside in a separate savings account which pays 4.00% interest compounded monthly. It deposits the first $28,000 today. If the company had wanted to deposit an equivalent lump sum today, how much would it have had to deposit? (Note that this is annuity due) $1,320,495.02 $1,470,961.78 $1,516,210.34 $1,588,534.50 $1,622,442.77
Answer #1

Answer:

Annuity A=$28,000

Interest rate=4% per annum

Per Month interest rate r=4%/12=0.33%

n=5*12=60 months

So PV of Annuity Due=A+A*(1-(1+r)^-(n-1))/r

PV of Annuity Due=28000+28000*(1-(1+0.33%)^(60-1))/0.33%

PV of Annuity Due=$1525441.84